OpenStudy (jakob):

Find the value of y when x equals 2. 2x-2y=22

OpenStudy (torvela):

once again fill in x for 2 to get: 4-2y=22

OpenStudy (ajspeller):

Substitute in the value of 2 for x 2(2)-2y=22 Then solve for y

OpenStudy (jakob):

Guys only TG sorry but he has helped me

OpenStudy (tgstudios):

just do the same process for all your questions @Jakob

OpenStudy (torvela):

you know what my bad for trying to help, but if your going to ask for help then not take it I will not even try.

OpenStudy (ajspeller):

lol .. okay dont have to tell me twice

OpenStudy (jakob):

Torvela sorry

OpenStudy (jakob):

So 2*2=4

OpenStudy (tgstudios):

4 - ? = 22

OpenStudy (jakob):

22-4=18

OpenStudy (jakob):

-18

OpenStudy (tgstudios):

YES! :D :D

OpenStudy (jakob):

That is the awnser

OpenStudy (tgstudios):

yup :D

OpenStudy (jakob):

Says it is wrong

OpenStudy (tgstudios):

ohhh... there is the 2 before y... *facepalm*, cut -18 in half

OpenStudy (jakob):

It is fine I cant miss anymore awnser

OpenStudy (jakob):

Alright new Q